34. In a factory, Ador and Joey make the same item. One day Ador made 39 more than three times what Joey made, and together they made 79. How many items did Joey produce that day?
Answer by ikleyn(52879)

.

    Ador     + Joey = 79


    (3x+39)  +  x   = 79


     3x      +  x   = 79 - 39


        4x          = 40


         x          = 40/4 = 10.


ANSWER.  Joey produced 10 items this day.

Solved, answered and explained without using words.